An engineer is going to redesign an ejection seat for an airplane. The seat was designed for pilots weighing between 140 lb and 201 lb. The new population of pilots has normally distributed weights with a mean of 147 lb and a standard deviation of 31.1 lb. a. If a pilot is randomly​ selected, find the probability that his weight is between 140 lb and 201 lb. The probability is approximately nothing. ​(Round to four decima
Answer #1

mean is 147 and sd is 31.1

P(140<z<201)= P((140-147)/31.1<z<(201-147)/31.1) = P(-0.225<z<1.736) =P(z<1.74)-(1-P(z<0.23))

from normal distribution table we get 0.9591-(1-0.5910)= 0.5501